A merchant wishes to mix coffee worth Php 18.00 per kilo with coffee worth Php 9.00 per kilo to get 90 kilos of mixture worth Php 12.00 per kilo. How many kilos of the Php 18.00 and 19.00 coffee will be needed?
Answer by Theo(13342)

x = amount of coffee worth 18 per kilo.
y = amount of coffee worth 9 per kilo

x + y = 90
18*x + 9*y = 12*90

from the first equation, you get y = 90 - x

substitute in the second equation to get:

18*x + 9*(90-x) = 12*90

simplify to get:

18*x + 810 - 9x = 1080

combine like terms and subtract 810 from both sides of the equation to get:

9x = 270

divide both sides of the equation by 9 to get:

x = 30

if x = 30, then y = 90-30 = 60.

your answer should be x = 30 kilo and y = 60 kilo.

30 * 18 = 540
60 * 9 = 540

540 + 540 = 1080 php total.

30 kilo at 18 php per kilo plus 60 kilo at 9 php per kilo will result in 90 kilos for a total of 1080 php.

1080 / 90 = 12 php per kilo.

your answer is confirmed as correct.